OpenStudy (anonymous):

Solve for x. x/7 = -3 x = -21 x = -10 x = 4 x = 21

OpenStudy (anonymous):

Okay, throw that 3 on the other side. x/7+3=? Then punch the x vales in for each equation -21/7 + 3 = 0 -10/7 + 3 = 11/7 and so on...

OpenStudy (tkhunny):

Have you considered simply multiplying by 7?

OpenStudy (anonymous):

Hey LoisCarmenDenominator, I think those bottom x answers are for multiple choice:-), but you had the right idea if that is what you would have been doing:-)

OpenStudy (anonymous):

x = -21 -_-